### v3

# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes (Experiment, Analysis & Benchmark Paper)

Abstract.
We present an in-depth analysis of data discovery in data lakes, focusing on table augmentation for given machine learning tasks. We analyze alternative methods used in the three main steps: retrieving joinable tables, merging information, and predicting with the resultant table. As data lakes, the paper uses YADL (Yet Another Data Lake) – a novel dataset we developed as a tool for benchmarking this data discovery task – and Open Data US, a well-referenced real data lake. Through systematic exploration on both lakes, our study outlines the importance of accurately retrieving join candidates and the efficiency of simple merging methods. We report new insights on the benefits of existing solutions and on their limitations, aiming at guiding future research in this space.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/rcap107/retrieve-merge-predict.
1. Introduction
In our data-rich era, data lakes –large, loosely structured corpora of raw data– have emerged as central resources in both academia and industry (Nargesian et al., 2019). Within data lakes, data discovery is crucial, particularly through the integration of multiple tables. This process has gained significant attention (Fan et al., 2023a). In this work, we focus on a specific usage of data lakes that requires capturing only a fraction of the tables: augmenting a base table for a machine learning (ML) task, such as in the following example scenario.
Alice, a data scientist, wants to predict the box office revenue of a movie: she has access to a table that contains information about movies (e.g., year of release, director, language, budget, …). While she can use the available data to train a model, finding more information on the subject would be beneficial; for example, joining the table about movies with a table on the cast of those movies might help, as some actors tend to appear in movies with higher revenue. Alice has access to some large repository of data, or to some search engine that she can query (Castelo et al., 2021; Google, 2023) to find additional tables. Her objective is therefore to find the tables that are most relevant to her task, to merge them with the original table and finally to use the improved table to build a model that predicts the box office revenue.
Figure 1. The pipeline. Given a base table, the three main steps (Retrieve, Merge, Predict) augment it with the information from the lake to improve the prediction performance.
As depicted in Figure 1, research on this problem is scattered across three main steps that involve four tasks: (1) retrieving the tables that are joinable with the original table based on shared attributes (Fernandez et al., 2019; Zhu et al., 2019, 2016; Castelo et al., 2021; Fernandez et al., 2018; Dong et al., 2023), (2) merging the information by executing the joins that most improve the performance of the subsequent ML model (Esmailoghli et al., 2021; Liu et al., 2022; Galhotra et al., 2023; Deng et al., 2017a), (3) aggregating results in cases of one-to-many or many-to-many joins (Chepurko et al., 2020; Kanter and Veeramachaneni, 2015), (4) predicting with the model on the resulting table (Prokhorenkova et al., 2018; Grinsztajn et al., 2022). For each task a variety of solutions have been proposed, yet each method is evaluated within disparate datasets and for the specific task it solves, rather than as the component of a full pipeline. In the absence of a unified benchmark to compare and evaluate these methods, it becomes hard for users like Alice to select the most suitable combination of methods for their specific use case.
Our study aims to promote a deeper understanding of discovery within data lakes for an analytic task. Our dataset, YADL (Yet Another Data Lake) makes such analysis reproducible and extensible. We leverage YADL to systematically explore the underlying mechanisms of end-to-end learning pipelines within data lakes.
YADL is constructed on the YAGO knowledge base (Mahdisoltani et al., 2014) to serve as a controlled environment for testing various methods by providing a defined vocabulary that includes entity names, thus alleviating issues stemming from uncertain matches and other ambiguities. The synthesis procedure accepts parameters to specify the shape (arity and cardinality) and degree of redundancy of the resulting tables: this allows to test how robust discovery and assembly systems are in scenarios with different degrees of complexity. YADL tables include both numerical and categorical features.
We investigate the impact of the main steps in the end-to-end process of learning from data lakes: generating potential augmentation candidates, integrating them with the base table, and subsequently evaluating the effectiveness of this assembled pipeline. We also provide a range of base tables to enable a comprehensive study.
This study has two main objectives: firstly, determine which tasks of the pipeline affect the prediction performance the most; secondly, develop an equitable playing field that facilitates a holistic and comparative assessment of alternative solutions.
Our pipeline enables the analysis of various methods in the four tasks. Our findings suggest that the retrieval of accurate join candidates is the most important task (improving performance by 4.9% in median relative to the worst retrieval method), outweighing the importance of the ML model used (4.1% difference in median). Results show that relying on Jaccard Containment (i.e., the fraction of entities of the table to augment found in a join candidate) as a criterion is effective for retrieving join candidates that improve downstream prediction performance. How join candidates are used to augment the base table affects the prediction performance, although the larger difference comes from joining multiple candidate tables at the same time rather than only selecting the best candidate (a difference of 4.2%). Also, simpler aggregation methods outpace their complex counterparts in terms of speed (being up to 4.5 times faster) and the larger cost of more intricate methods does not yield proportionate gains in prediction performance.
To validate the generality of our findings, we contrast the insights obtained from the analysis carried out via YADL with those obtained from Open Data US, a well-referenced real data lake (Nargesian et al., 2018; Galhotra et al., 2023), over 8 input base tables. This comparative approach ensures that YADL is representative enough to be a benchmark.
After outlining the problem setting and previous work in Section 2, we share our main contributions:
  1. (1)
    We develop YADL: a novel benchmarking data lake that allows to test retrieval and augmentation techniques in a controlled environment (Section 3). YADL, the base tables, and the pipeline are available and easily extendable to spur further research.
  2. (2)
    We implement a full prototype pipeline for augmenting tables from data lakes (Section 4). We then measure the prediction performance, execution time, and RAM usage of each method.
  3. (3)
    We conduct an experimental study on 8 base tables over various data lakes, retrieval techniques, join selection methods, aggregation solutions, and ML models (Section 5). Results provide insights on the steps of pipeline learning from data lakes.
We conclude the paper with a discussion of future research directions in Section 6.
2. Problem setting and related work
Problem setting. Consider a user training a ML model to predict some quantity. The corresponding quantity appears in the training data as a target column $Y$ of a base table $T$. Assume our user has also access to a large collection of tables (a data lake) $D=\{T_{1},T_{2},...,T_{m}\}$, some of which may contain additional information that can enrich the base table $T$. The data lake may be publicly available (Castelo et al., 2021; Mattmann et al., 2018; Srinivas et al., 2023; Hulsebos et al., 2023a; Nargesian et al., 2018) or private (e.g., corporate data). Each of these tables $T_{k}$ is a bi-dimensional collection of data organized in columns $C^{i}_{k}\in T_{k}$ that may include both categorical (names, codes etc.) and numerical data (price, revenue, tax rates etc.). While cross-table metadata such as foreign keys is not available in this setting, joining $T$ with some of the tables in $D$ would be beneficial for the target prediction task.
Given $T$ and $D$, a table $T_{k}\in D$ is considered to be joinable, or a join candidate, if at least one of the columns $C^{i}_{k}\in T_{k}$ has a non-empty intersection with one of the columns in table $T$: if ${}\exists Q\in T,~{}\exists C^{i}_{k}\in T_{k}~{}|~{}Q\cap C^{i}_{k}\neq\emptyset$.
The user is interested in optimizing the performance of the ML model on a collection of columns (or “features”) $X$ according to the quality of the prediction for the target $Y$ (e.g., the movie revenue). Columns in $X$ may come either from $T$, or from joined tables in $D$.
The different steps, and corresponding literature. In such a scenario, the user is likely to go over the following main steps: Join Candidate Retrieval, Join Candidate Merging (which includes the Join Selection and Aggregation tasks), and Prediction. Depending on the specific scenario, some of the tasks may be executed in a different order or not at all. In the following, we will drop Join from the names when clear from the context.
While we focus on the joinability problem, a related concept is that of table unionability, i.e., identifying which tables may be “appended” to $T$ to increase the number of rows without increasing the number of features (columns) (Khatiwada et al., 2023a; Fan et al., 2023b; Nargesian et al., 2018).
Finding the join candidates
Given a base table $T$ and a data lake $D$, the Candidate Retrieval task consists in discovering join candidates (“Retrieval” in Fig. 1). This task looks for tables that can be considered as “candidate joins” for the given base table. Depending on the scenario, different methods have been developed: some systems (Castelo et al., 2021; Mattmann et al., 2018) work as “search engines” of sort by querying an open data internal database; other methods (Zhu et al., 2019; Fernandez et al., 2019; Zhu et al., 2016; Fernandez et al., 2018; Khatiwada et al., 2022, 2023b) are instead instantiated “locally” on data that the user has direct access to. While integrating data is not a central focus of the ML literature, augmenting features via joins is recognized as key to ML (Paleyes et al., 2022; Kumar et al., 2016). Still, the blind addition of features may lead to diminishing or negative returns, which led to the development of systems that seek to integrate only correlated features (Esmailoghli et al., 2021; Santos et al., 2021). Alternative systems seek to augment $T$ by synthesising new features (Kanter and Veeramachaneni, 2015; Cvetkov-Iliev et al., 2023; Zhao and Castro Fernandez, 2022). An orthogonal research direction is concerned with the problem of search and augmentation while maintaining privacy (Huang et al., 2023a, b).
Though implementations may vary, retrieval methods share some similarities: 1) they involve an offline phase to either index the data lake (Zhu et al., 2016; Fernandez et al., 2019), or train a model (Fan et al., 2023b; Dong et al., 2023), and 2) they rely on similarity metrics across columns to select candidate joins. Often, this metric is Jaccard Containment (JC), defined as
(1) Jaccard Containment := $\frac{|Q\cap C_{k}^{i}|}{|Q|},$
where $Q\in T$ is a query column, $C_{k}^{i}\in D$ is a candidate column in table $T_{k}$, $|Q\cap C_{k}^{i}|$ is the cardinality of the intersection between the two sets, and $|Q|$ is the cardinality of the query set itself. Intuitively, if this ratio is high, then a large fraction of $Q$ is found in $C_{k}^{i}$, suggesting that the two columns should be joined. Exhaustively measuring the JC is expensive. Thus approximate methods have been developed for fast estimation of the containment (Zhu et al., 2016; Fernandez et al., 2019, 2018). Beyond Jaccard Containment, other metrics have been used, such as top-$k$ set similarity (Zhu et al., 2019), combinations of JC and embeddings-based similarities (Fan et al., 2023b), or other embeddings-based metrics (Cong et al., 2022; Habibi et al., 2020), as approximate nearest neighbor search (Dong et al., 2023) on language models. Methods based on Jaccard Similarity are vulnerable to typos, which led to the development of methods that perform fuzzy or semantic matching such as (Fan et al., 2023b; Mundra et al., 2023; Deng et al., 2017b); in this work, we focus on scenarios that do not involve fuzzy matching.
Retrieval methods are typically designed to be used on large data lakes (potentially, up to millions of tables (Zhu et al., 2019; Fan et al., 2023b)) and be as exhaustive as possible to maximize recall. However, three issues arise. First, these methods do not assess the relevance of join candidates to the downstream task. While a large containment value may indicate that a join can be done, it provides no guarantee that this join is actually useful (Kumar et al., 2016). Second, the number of candidate joins could become too large for practical use. Manually identifying the best candidates is time-consuming, and performing all joins might be too expensive in terms of time/memory constraints (Santos et al., 2021). A user-defined threshold on the containment can filter the most promising joins, but deciding the correct threshold is problematic. Alternatively, it is possible to select only the top-$k$ candidates by containment. Third, Jaccard Containment does not take into account the cardinality of a column: in an extreme case, if a column $Q\in T$ contains only a single value, it would have perfect overlap with any column $C_{k}^{j}\in D$ that contains the same value. While this behavior may be desirable for some specific retrieval tasks, it further expands the number of candidates with high containment. These limitations have led to the development of methods tackling the Join Selection task. As executing joins can be expensive, different methods have been proposed to reduce this cost, from avoiding the join altogether (Kumar et al., 2016; Shah et al., 2017), to joining over sketches or coresets of the data (Wang et al., 2022; Santos et al., 2021). A more straightforward approach consists in testing the candidate joins and select only those that improve the prediction performance (Chepurko et al., 2020; Gong et al., 2023; Galhotra et al., 2023; Dong and Oyamada, 2022).
Merging the candidates
After the set of join candidates has been retrieved, they get merged with the base table to augment it with their columns (Merge in Fig. 1).
The Join Selection task addresses the problem of identifying a subset of candidate joins that maximizes the prediction performance over a downstream task. Candidate selection is performed by profiling candidate joins according to various parameters (Galhotra et al., 2023; Flores et al., 2021), removing joins that are not useful through the application of rules (Kumar et al., 2016; Shah et al., 2017), or by executing each candidate join to find those that bring benefit (Esmailoghli et al., 2021; Chepurko et al., 2020; Liu et al., 2022; Galhotra et al., 2023). In this work, we focus on the latter strategy.
It is therefore necessary to address the difficulties that come from joining base table and candidates. We consider ML models that involve a specific set of samples (rows) with their features. Augmentation operations in this scenario enrich said features while keeping the original set of samples constants, thus requiring a left join. A complication with left joins is that, for any un-joined row in the left table, null values will be added in the new columns. Alternative methods based on Full Disjunction (Galindo-Legaria, 1994; Khatiwada et al., 2022) add irrelevant samples that should then be dropped.
Note that, while we split the Candidate Join Retrieval and Selection steps for clarity, there is no hard separation between them. Indeed, some methods cover multiple steps (Gong et al., 2023; Huang et al., 2023c).
It is not always possible to limit joins to one-to-one relations; often, we join on one-to-many or many-to-many relations. For example, to join a table about movies with one that contains movie ratings on column “movie title”, every movie with more than one rating is involved in a one-to-many relation, thus the content of all the rows with such relations gets duplicated. Sample duplication is problematic when the downstream task involves a ML or statistical analysis method, as such models assume that each row correspond to one sample with its features. Thus, the Aggregation task bridges the result of the join with the downstream methods, combining the information contained from a potentially large number of rows into one. More precisely, given a tuple $t$ in the base table $T$ that needs to be joined with $n$ tuples from a binary table ($A$, $B$) over $A$, the goal is to augment $t$ with attribute $B$ by selecting one value that represents the information from the $n$ joining tuples. How to select a representative value for the new attribute reminds traditional data integration problems such as truth discovery (Dong et al., 2009) or data fusion (Bleiholder and Naumann, 2009). In this spirit, Deep Feature Synthesis (DFS) (Kanter and Veeramachaneni, 2015) takes a set of tables and a join plan to recursively aggregate replicated instances using functions such as average, median, and mode.
Learning on the augmented data
Finally, the integrated table is used to train a model to predict a target variable through simple methods such as linear regression/classification (Pedregosa et al., 2011), or more complex ones such as CatBoost (Prokhorenkova et al., 2018). Though we present the prediction task as distinct from the prior ones, an optimized pipeline may implement learning at intermediate sections of the pipeline to ensure the selection of the most appropriate candidate tables(Galhotra et al., 2023; Huang et al., 2023c).
3. Building Yet Another Data Lake
We detail the construction of YADL (Yet Another Data Lake), a synthetic data lake built starting from the YAGO (Suchanek et al., 2007; Mahdisoltani et al., 2014) knowledge base (KB). Our goal is a scalable and reliable data lake that allows users to evaluate the main steps of the pipeline in Fig. 1, while avoiding some of the confounding factors that come from working with unvetted data, such as typos, inconsistent schemas and data format, and other sources of noise. These challenges can easily be added to YADL, e.g., by generating typos in table entries.
3.1. YAGO: the source of the original data
YAGO (Suchanek et al., 2007) is a KB composed of RDF triplets; each triplet has a subject connected to an object through a predicate (or relation). Subjects and objects are considered to be entities, e.g., in the triplet “Paris, locatedIn, France”, “Paris” is the subject entity, “France” is the object entity, and “locatedIn” is the predicate that connects them. The object may also be a lexical value such as the “population density”. Each entity belongs to a set of classes (or types), arranged in a taxonomy; “Paris” belongs to the class “City”, subclass of “Populated place”, itself subclass of “Geographical locations”. We use YAGO 3.0.3 (Mahdisoltani et al., 2014), which is updated to 2022.
3.2. From knowledge base to relational tables
Information in data lakes is typically stored as tuples in tables, rather than the triplet format used in YAGO. For this reason, we construct YADL by first reshaping the triplets into binary tables. Wide-form relational tables can then be constructed by joining binary tables on the column that contains the triplet subjects. We assemble two types of YADL variants (Binary and Wordnet-based), which differ in the number of tables and in their shape (i.e., the number of rows and columns of each table).
Binary tables
A binary table is generated for every predicate in YAGO, e.g., “hasCapital”. YAGO contains 70 predicates, which result is a small data lake where some of the tables have millions of rows (e.g., “isLocatedIn”) and some have few (e.g., “hasTLD”). Each table contains an attribute named “subject” and another attribute named as the actual predicate. For example, the triplet “France - hasCapital - Paris” leads to a table “hasCapital” with columns “subject” and “hasCapital”; finally, a row “France, Paris” is added to the table. The same process is applied to all 70 relations and their triplets.
Table 1. “City” seed table built by flattening predicates.
Subject locatedIn_1 locatedIn_2 owns_1 density
Paris France Europe Tour Eiffel null
Rome Italy null Olympic Velodrome 2236
Table 2. “City” seed table built by left-joining binary tables.
Subject LocatedIn Owns Density
Paris France Tour Eiffel null
Paris Europe Tour Eiffel null
Rome Italy Olympic Velodrome 2236
Wordnet-based tables
To reflect that tables typically have more than two columns, we develop a second variant of YADL where tables have a larger number of columns. To this goal, we leverage the Wordnet (Miller, 1994) classes (e.g., “Person”, “Company”, “Artist”) to which entities belong. We use all Wordnet classes (a total of 1015) to create our seed tables.
For each class, a new table is generated following the example in Table 1. Initially, the seed table includes solely the “Subject” column (e.g., the “City” table includes only values “Paris,” “Rome,” etc.). Subsequently, we join every relation associated with subjects in the table: in Table 1, the subjects with type “City” are joined with relations “locatedIn”, “owns”, “hasPopDensity”.
Transforming triplet tables into wide format tables leads to tables with null values because not all entities of a given type have the same relations. For example, column “hasPopDensity” contains null values for subject “Paris” in Table 1. The fraction of missing values in the transformed tables is reported in Table 3.
One-to-many relations
A subject may be linked to many objects through the same predicate, e.g., “Paris” is linked to two objects (“France”, “Europe”) via the predicate “locatedIn”. We address the issue in three ways. For Binary, triplets that share the same subject and predicate are transformed into tuples with the same value in the “subject” column and different values in the predicate column, e.g., the binary table has columns “Subject” and “isLocatedIn” and it contains tuples (Paris, France) and (Paris, Europe). For “Wordnet” seed tables, we build two variants. In the first case (which we dub YADL Base in section 5), we flatten the group of objects by creating new columns, thus moving them on the same tuple rather than splitting them (e.g., in Table 1, column “locatedIn” is flattened over “locatedIn_1” and “locatedIn_2”); in the second variant (YADL 10k and YADL 50k), we create the new columns by executing a left join between the subject and each binary table. One-to-many matches (such as with “Paris” and “France”, and “Paris” and “France”) are replicated with each left join; ordinarily, this would lead to an explosion in the number of rows: we limit the issue by selecting only the first two entries for each predicate (as shown in Table 2).
As a result of handling these relations differently in each version of YADL, the aggregation step in the pipeline affects them in different ways, thus exposing different problems.
Sub-tables
We create supplementary tables derived from the Wordnet seed tables to augment the table count in the data lake with two different methods. In real-life lakes, numerous versions or variants of the same table are common, encompassing collections of tables that undergo slight modifications over time (Halevy et al., 2016). This redundancy poses a challenge for information retrieval methods in distinguishing between pertinent and extraneous tables. Hence, it constitutes a crucial aspect in a benchmark data lake like YADL, which incorporates it heavily in the generation of sub-tables.
For YADL Base, we generate all combinations of arity 2 and 3 for each seed table. For Table 1, these combinations would include “(isLocatedIn_1, owns_1)”, “(owns_1, popDensity)”, and “(isLocatedIn_1, owns_1, popDensity)”. Each sub-table is then built by projecting the seed table over the generated combination. We exclude all sub-tables with fewer than 100 rows. Rows containing null values are retained if the column combination includes at least one non-null value. For YADL 10k and 50k, we drop all seed tables whose arity is lower than a minimum arity $A$; then, for each surviving seed table $T$, we generate $N$ sub-tables by sampling a random subset of columns of size $[A-2,A]$ and projecting onto $T$; each parameter can be tweaked to modify the size and number of resulting sub-tables. Row-wise redundancy is provided by randomly sub-sampling a fraction $p$ of each resulting sub-table $n_{s}$ more times, thus replicating a fraction of the samples while keeping the set of columns fixed. We construct YADL 10k and YADL 50k by setting $N$ to 10 and 50 respectively, $A=8$, $p=0.7$, and $n_{s}=2$.
Data lake statistics are reported in Table 3, while Figure 2 reports the measured containment for the base tables we use in our experiments..  [^1] The data lakes are available at https://zenodo.org/doi/10.5281/zenodo.10600047, while the code to prepare YADL is at https://github.com/rcap107/YADL.
Figure 2. Exact containment measured over the top 100 join candidates for each base table (represented by a colored line) on each data lake. The x-axis reports the rank, rank 30 is the retrieval cutoff we use in our experiments.
4. Implementation of the pipeline
We now discuss the steps of the pipeline in Figure 1. For the Retrieve step, we describe the advantages and disadvantages of each candidate retrieval method; for the Merge step, we propose different join selectors and aggregation methods, and we explain how these two parts are intertwined; finally, for the Predict step, we go over the ML methods used to test the prediction performance.
4.1. Retrieving the candidates
We consider retrieval methods that work by taking a query column and return a – possibly ranked – list of candidates.
Retrieval methods usually involve offline preparation steps such as building sketches (Zhu et al., 2023; Fernandez et al., 2019), indices (Zhu et al., 2019), creating embeddings (Fan et al., 2023b), or measuring the Jaccard Containment of each column; preparation includes persisting the data structures for future use. Afterwards, candidate joins can be obtained by querying the retrieval method.
For retrieval strategies estimating Jaccard containment (Equation 1), a candidate column is a good join if it contains a large fraction of the entities of the query column (Section 5.2 provides an empirical justification). We consider two complementary strategies, an exact and a stochastic approximation, as well as a “hybrid” approach combining them. Starmie (Fan et al., 2023b) is our reference baseline.
The rest of the pipeline is transparent to the retrieval method: given a list of join candidates, it will test each candidate regardless of how the list was constructed. For our experiments, we reduce the number of candidates to the top-30, as we observe that almost all candidates with high containment are within this limit (Fig. 2).
Exact Matching
We compute Exact Matching (Exact) by measuring the exact Jaccard containment for each pair (query column, candidate column) in the data lake. This can be implemented efficiently by first building a “vocabulary” on the query column, and then scanning the whole data lake to compute containment. Candidates are ranked by highest containment, then we select the “top-$k$” candidates in the pool. An advantage of this method is that computing the containment is an embarrassingly parallel operation.
The main drawback of Exact Matching is the computation has a cost that depends directly on the size of the data lake and the tables therein, and this operation must be repeated for every new query column. As we show in Fig. 10, the cost for multiple query columns quickly adds up. Another drawback is that containment is less reliable if the data lake features a lot of redundancy (e.g., YADL 50k in Fig. 2), since unrelated tables may still feature high containment.
MinHashLSHEnsemble (MinHash)
MinHash (Zhu et al., 2016) relies on Locality Sensitive Hashing (LSH) to build an index subject to a user-set minimum containment threshold. At query time, all candidate columns with estimated containment larger than the threshold are returned. For consistency with the other methods, we select $k$ candidates from this pool.
The main drawback of MinHash is its lack of a candidate ranking, which means that selecting a subset of the candidates does not guarantee that the candidates will have high containment (Figure 8a). Moreover, since MinHash returns an approximate result, it may happen that candidates that have a measured containment lower than the threshold are proposed: indeed, when the MinHash query results are compared with the results obtained from Exact Matching, MinHash produces a large fraction of False Positives. Finally, while the index can be updated when new data is added to the data lake: performance may deteriorate after a certain level of updates, especially depending on the skew of the new data.
Hybrid MinHash
To address Minhash’s lack of a candidate ranking, we propose Hybrid MinHash. Hybrid MinHash leverages the prebuilt index from MinHash to find candidates for a given query column, then uses Exact Matching to rank all candidates. The “top-$k$” candidates are then selected from the ranked list. By measuring the containment over the MinHash query result, the pool of candidates is reduced compared to testing all columns in the data lake (like with Exact Matching).
Just as it takes some of the strong suits of both MinHash and Exact, Hybrid MinHash also retains some of their disadvantages: 1) The MinHash index will degrade each time it is updated. 2) As the first filtering relies on the MinHash results, False Positives add needless cost and any candidate missed by MinHash is lost. 3) Re-ranking candidates by calculating the exact containment increases the query time substantially with respect to MinHash.
Starmie
We compare the simple methods described so far against Starmie (Fan et al., 2023b), a SOTA system for table unionability and join discovery. Starmie employs a contrastive learning method to train column encoders from pre-trained language models to capture contextual information within tables. Starmie performs candidate retrieval by building embeddings for each column in the data lake and for the query column; it then ranks candidates with a similarity score combining embedding cosine similarity and Jaccard Containment.
4.2. Merging base table and join candidates
In this step, the candidates that were retrieved in the previous step are combined into a new integrated table that joins the information in the base table $T$ with additional information from augmentation tables. As we highlight in Fig. 1, the merge step combines two operations that cannot be executed independently of each other: join selection, and aggregation.
Aggregation
When joining tables for a downstream machine learning task, one-to-many relationships must be aggregated to avoid replicating samples in the base table which would modify the initial data sampling.
Figure 3. Example of how a left join would duplicate rows from the base table.
Following the example in Fig. 3, a join on “Title” leads to the duplication of the two first rows in the base table, as both title values appear in two rows of the candidate table. How should we aggregate the rows to obtain the best prediction over the column “Revenue”?
We test three different aggregation strategies:
  1. (1)
    First drops all rows where the values belonging to the base table are duplicated, with the exception of the first copy.
  2. (2)
    Mean replaces all duplicated numerical (categorical) values by the mean (mode) of all values for that attribute.
  3. (3)
    DFS (Deep Feature Synthesis (Kanter and Veeramachaneni, 2015)) greedily adds new features by performing different aggregations (mean, median, count, standard deviation, etc.) over each column in the table.
Join selection
We implement and test five different join selectors. All selectors follow the fit-predict paradigm proposed by scikit-learn (Pedregosa et al., 2011), which makes extending the pipeline with new selectors very simple. All experimental runs are executed over 10 cross-validation splits: the test fraction is 0.2 in this outer cross-validation split. Each selector receives as input a pool of “$K$” candidates from a given retrieval method on a given data lake, the train and test splits of the base table, and the aggregation method to use. The train split is further split into a training (0.8) and validation set (0.2). The model is trained on the final integrated table (i.e., the base table after joining it with the candidates) produced by the selector of this training split, then the result is scored with the test split.
The reference selector is No Join, where the ML model is trained directly on the training split of the base table.
Highest Containment Join measures the exact containment of all of the candidates, then re-ranks them by that. The candidate with the highest containment is selected and joined; ties are broken by taking one candidate at random.
Best Single Join iterates over each candidate one at a time, performs the aggregation and join, then trains an ML model on the joined table. The candidate with the highest model performance on the validation split is selected and joined.
Top-$k$ Full Join aggregates and joins the first $k$ candidates as ranked by a retrieval method, without performing any re-ranking.
Full Join aggregates and joins all candidates from the pool.
Stepwise Greedy Join aims to add more information than what is available in a single table, while reducing the amount of noise added by joining irrelevant tables. The baseline performance is measured by training a model on the base table’s training split and validating on the validation split; candidates are re-ranked by their containment before iterating over each of them. During each iteration, the candidate is aggregated and merged separately with each of the two splits of the current table; a new model is trained on the training split and validated on the validation split. If the performance improves compared to the previous iteration, the newly joined table becomes the new current table, otherwise the candidate is discarded.
Highest Containment and Best Single Join produce smaller integrated tables as they only join one candidate, rather than all potential candidates like Full Join and Stepwise Greedy Join: we can therefore classify the two pairs as single-table selectors and multi-table selectors respectively. Top-$k$ Full Join can be assigned to either class depending on the value of $k$.
In all selectors that involve a join, aggregation is carried out prior to executing the join itself.  [^2] In our pipeline, aggregation is carried out before executing any join by grouping the “right table” by the join key, then applying one of the aggregation functions described above. This is to avoid materializing large joins, which have a huge cost in memory and time. Due to how aggregation is carried out, the final result is the same before and after the join, so executing it before materializing the joined table is more efficient.
4.3. Predicting with a ML model
The learning step is performed on the final integrated table on the training split returned by the join selector and evaluated on the integrated table on the test split. We evaluate two different methods: a simple linear regressor/classifier (Pedregosa et al., 2011) and the more complex method CatBoost (Prokhorenkova et al., 2018). We choose CatBoost as our “complex and effective” method because gradient boosting methods are effective at working with tables (Grinsztajn et al., 2022), and CatBoost in particular is good with categorical data, which is common in our scenario.
5. Experimental study
For our experimental campaign, we test the different sections of the pipeline over three main dimensions: prediction performance, execution time, and RAM usage. Focusing on multiple dimensions allows us to have a view of the different trade-offs across methods.
Table 3. Aggregate statistics for the tables in the data lakes, along with the size on disk of the data structures used by each retrieval methods. We report the mean (Avg) for each statistic. “N. cols” and “C. cols” refer to numerical and categorical columns respectively.
YADL YADL OpenData
Binary Base 10k 50k US
N. tables 70 30072 10059 47223 5591
Size (MB) 628.9 10051.3 5557.7 27718.1 4062.4
Tot. rows 2.01e+07 6.72e+08 2.42e+08 1.20e+09 9.57e+07
Tot. cols 1.40e+02 9.52e+04 1.27e+05 6.24e+05 1.33e+05
Avg rows 287134.3 22343.9 24014.8 25430.2 17124.5
Avg cols 2.00 3.17 12.63 13.21 23.86
Avg N. cols 0.30 0.39 3.54 3.59 11.10
Avg C. cols 1.70 2.78 9.08 9.61 12.76
Avg nulls 0.00 0.31 0.62 0.64 0.09
Data lakes
We use four YADL variants (Binary, Base, 10k, 50k) (as detailed in Section 3) and Open Data US, a data lake employed in the literature (Galhotra et al., 2023; Zhu et al., 2019, 2016); statistics are reported in Table 3.
Table 4. Statistics for the base tables.
Base Table Target # Num. Att. # Cat. Att. # Rows
Company
Employees
Number of
Employees
2 7 3109
Housing
Prices
House
Price
3 7 22250
Movies
Movie
Revenue
8 10 7397
Schools
School
Class
4 3 1774
2021 US
Accidents
Accidents
by County
1 3 14850
2017-2022
US Accidents
Accidents
by County
2 3 80296
US County
Population
County
Population
1 1 3059
2020 US
Election Results
Vote %
by County
3 6 22093
Base tables
We evaluate six tables from sources that are not related to the lakes: Company Employees, Movies, US Elections, 2021 US Accidents, 2017-2022 US Accidents, and Housing Prices.
We also include one “internal” table derived for each lake: US County Population (from YADL), and Schools (from Open Data US). We include internal tables as those have been used in several previous works for evaluation (Galhotra et al., 2023; Chepurko et al., 2020).
For all datasets, the query column values must be matched with the entities in YADL using semantic annotation solutions (Huynh et al., 2022; Nguyen et al., 2021). In our experiments, we have done the match manually to remove the noise from this task. Also, query columns are chosen based on what a user may reasonably consider as “key” (e.g., the movie title).
To reflect the experience of a data scientist that needs to construct a meaningful table starting from a data lake, and to highlight the effect of joins on the downstream task, we run experiments on a depleted version of the tables, i.e., the input tables include only the primary key column and the target column.
Datasets’ statistics are reported in Table 4.
Compute resources
Experiments were executed on a single server with 72 Intel Xeon Gold 6140 CPUs @ 2.63GHz and 376GB of RAM.
Implementation details
The implementation of the pipeline is in Python; Exact Matching, aggregation and join operations are implemented using Polars (Vink et al., 2024) as backend, while the ML models are implemented using Scikit-learn (Pedregosa et al., 2011) and CatBoost (Prokhorenkova et al., 2018). To improve run consistency on the shared server, we limit the number of Polars and Catboost threads to 32.
We rely on the Python implementation of MinHash provided by the Datasketch package (Zhu et al., 2023). For Starmie, we optimize the original implementation so it can scale to work in our environment.  [^3] The repository we used is available at https://github.com/rcap107/starmie. We use the official DFS package. Data structures are stored by persisting on disk the ensembles for MinHash, the candidate ranking for each query column for Exact Matching, and the model checkpoints for Starmie; Hybrid MinHash relies on the MinHash data structures to work, so no additional storage is required for it. Table 6 reports the size on disk of the data structures used for each method.
Retrieval We use a containment threshold of 0.2 for the preparation of the MinHash index, and clamp the number of candidates returned by each retrieval method (Exact Matching, MinHash, Hybrid MinHash, and Starmie) to 30. These values were chosen to balance execution time and expected number of candidates given the distribution of containment encountered in the different data lakes (Fig. 2). For Starmie, we use the default parameters defined in the original repository.
Selection We fix the number of Stepwise Greedy Join iterations to 30: this number is consistent with the number of candidates that are provided in the retrieval step. For Top-$k$ Full Join, we set $k=1$ to compare the performance of the ranking provided by each retrieval method with Highest Containment. None of the other join selectors have feature parameters to tweak.
Learning We fix the number of CatBoost iterations to 300; we stop training the model 10 iterations after the optimal metric has been detected; we set the L2 regularization coefficient to 0.01.
5.1. Retrieval is the most impactful pipeline step
Our first goal is to pinpoint which steps of the pipeline have a significant impact on the studied dimensions (if there are any): optimizing these influential sections can yield the greater benefits.
We do so by separating the variance that is due to that specific task from that due to the dataset and the other tasks for each pipeline task (retrieval, selection, aggregation, learning). This is achieved by aggregating variables other than the one of interest to find the “average method”, which becomes the reference point (either 0% prediction, or 1x run time) in each plot. For instance, when analyzing the impact of the retrieval step, results are grouped by join selection model, aggregation method, and ML model before averaging prediction performance and run time. Then, the results obtained with each value in the variable of interest (e.g., Exact Matching, MinHash and Hybrid MinHash) are plotted as a difference from the average method for that variable (the “average retrieval method”). When dealing with binary variables, such as the ML model, the “best” method is identified using the median prediction score score, then we plot the difference from the alternative. We report the difference in prediction score ($R^{2}$ and AUC) and relative execution time with respect to the average method for retrieval (Fig. 4), selection (Fig. 5a), aggregation (Fig. 5b), and learning (Fig. 5c). Individual folds are reported as dots; color palettes depend on the data lake.
Figure 4. Comparison between different retrieval methods. The left side (retrieval) reports the computational cost of each method; the right side (prediction pipeline) reports the prediction performance and time required to run the prediction pipeline. All results are reported with respect to the “average method”. The median difference is reported on the right side of each plot.
a. Join selection method
b. Aggregation method (only join selection with Best Single Join and Highest Containment Join)
c. Supervised learner
Figure 5. Experimental results across all data lakes, comparing the performance difference between evaluated methods in different steps of the pipeline. The median difference is reported on the right of each plot. (a) compares join selectors, (b) aggregation methods, and (c) ML models. In all cases, results are relative to the “average method”. Starmie results are not reported in this plot.
Figure 6. Relative prediction performance (median on the right) obtained selecting the (top-$1$) candidate according to the ranking provided by the retrieval methods.
Retrieval
Observations on the retrieval methods are shown across two dimensions: the resource cost incurred in building the retrieval structures, and the effect of each retrieval method on the prediction pipeline. Fig. 4 (left) highlights the huge difference in both peak RAM (more than 200x in median) and training time (about 13x in median) required to prepare Starmie, especially when compared to the other methods. The very large RAM cost is particularly problematic for Starmie because it prevents the method from running on most GPUs when working with larger data lakes. Exact Matching has a very low peak RAM usage because containment can be measured by loading in memory a single candidate table at a time, and there is no need for sizeable data structures (like with MinHash), or feature vectors (Starmie). More detail on the recorded peak RAM is reported in Table 6.
Regarding the prediction performance (right Fig. 4), we observe that Exact Matching shows the best overall prediction performance (2.25% gain in median across all experiments), while MinHash performs substantially worse (-2.60%); Hybrid MinHash shows a marked improvement compared to MinHash. The prediction performance of Starmie is similar to that of Exact Matching, which makes sense as the similarity metric used by Starmie combines Jaccard similarity with the cosine similarity of column embeddings. MinHash is faster than the others, despite the fact that the number of candidate joins $k$ is 30 for all methods. This is due to the fact that, on average, the candidates retrieved by MinHash have a much lower containment than those proposed by the other two methods (Fig. 8a): this results in a shorter join and train time due to the smaller amount of data that must be moved and used for fitting the models. The difference between Hybrid MinHash and the two methods based on Jaccard similarity is likely due to MinHash does not having perfect recall: some of the higher containment joins are replaced by candidates with lower containment. Overall, the two methods based on precise ranking (Exact Matching and Starmie) outperform the methods based in approximate matching (MinHash and Hybrid MinHash), suggesting that the computational cost incurred in measuring exact containment does result in better prediction performance.
Selection
The choice of selector shows a noticeable effect when moving from single-table selectors (Highest Containment and Best Single Join) to multi-table selectors (Stepwise Greedy and Full Join), bringing a benefit close to 4.3% in median (Fig. 5a).
Stepwise Greedy Join is an outlier in how much slower it is compared to all other methods. This is not surprising, since this selector executes a join and trains a model in each iteration, then re-trains the model at the end of the fit step. This selector is expected to yield better results than Full Join by avoiding candidate joins that do not bring substantial benefit(only 0.04% in median): this makes the much longer execution time even more concerning.
Highest Containment is significantly faster than all other solutions because it only re-ranks candidates before joining the best.
The difference in performance between the two single-table selectors (Best Single Join and Highest Containment) is likely due to the fact that, while the Jaccard Containment is a good indicator of a potentially good join, it is not sufficient for selecting the best candidate. In fact, multiple candidates may have the same exact containment value, leading to ties. Highest Containment breaks ties at random, thus it may select tables that are not related to the base table. Data lakes that feature a large amount of redundancy like YADL 50k, or YADL Base (Figure 2) are more affected by this problem. The significant difference between single-table selectors and multi-table selectors is explained by the learning model benefiting from an increased set of features: merging more than one table inherently results in a richer feature set compared to the scenario where only a single table is joined.
An important observation is that all selectors rely on the candidates proposed by a retrieval method: if these candidates have poor quality, the selectors cannot compensate for that.
Figure 6 reports the prediction performance across different data lakes obtained by joining each base table on the top-$1$ candidate according to each retrieval method’s internal ranking (here, Exact is equivalent to top-$1$ Highest Containment Join). Unsurprisingly, MinHash is far worse than all other methods (up to 6.5% in median) due to the lack of an internal ranking; Hybrid MinHash performs much better than base MinHash, however the imperfect retrieval performance penalizes it. Starmie and Exact Matching have similar performance, though it appears that the first candidate retrieved by Starmie is not as good as what is selected by Exact Matching.
Aggregation
Among the pipeline steps under evaluation, aggregation exhibits the least variance in prediction performance between methods (a maximum of 0.65% in median). DFS does improve the performance, however at the cost of a much longer execution time (up to 4.5x longer than simpler methods). It also creates new features greedily, which may lead to memory issues (Cvetkov-Iliev et al., 2023).
Experiments reported in Fig. 5b were obtained using only a subset of the datasets (Company Employees, Movie Revenue, Schools, US Accidents 2021, US County Population), and include only the single table selectors because DFS ran into scalability issues with Full Join and Stepwise Greedy. Combining DFS with the Stepwise Greedy selector would skew the time difference plot even further.
Learning
The comparison between learning methods yields predictable results: CatBoost outperforms the linear model (4.14% in median); it is however about 60% slower in median, likely due to the fact that some runs do not converge and early stopping does not help. This outcome is not very surprising, since CatBoost is a complex, state-of-the-art model. Due to the clear trade-offs in this step, the user should make their model choice based on these considerations and their requirements.
Figure 7. Comparison in prediction performance for the different data lakes.
Different Lakes
Figure 7 reports aggregated absolute gain from the un-joined base tables on the different data lakes to highlight the potential gains in prediction performance that can be achieved by augmenting them with candidate joins. The figure also shows the risk involved in joining sub-optimal candidates (shown by the negative prediction performance). Open Data US shows the worst prediction results overall: these results can be explained by the inherent “dirtiness” of the data lake, especially when compared to YADL: Open Data has a lower average containment compared to YADL (Figures 8a and 2). On the other hand, the good results can be attributed to the “Schools” internal dataset, which achieved perfect classification performance in most cases and skewed the results in that direction (also visible in Figure 8b, where highest containment leads to best performance). YADL’s internal table (US County Population) exhibits a similar behavior to that of Schools. Such behavior can be explained by the fact that internal tables, i.e., tables that are sampled from the data lake itself, tend to have copies in the data lake, which are likely to contain information that is correlated with the prediction task. This highlights the specific behavior of internal tables, which are routinely used in the experimental campaigns in the literature. The different prediction performance obtained with each YADL variant highlights the difference in the internal structure of each data lake. The reason why YADL 10k and 50k underperform wrt YADL Base might be found in the larger fraction of null values (Table 3) and, in the case of YADL 50k, the larger degree of redundancy (Fig. 2).
An advantage of YADL is that it is built based on a KB that contains “general knowledge” notions: as a result, external tables are more likely to find matches in YADL than in Open Data US, whose tables contain mostly US-based demographic data.
5.2. Containment affects the entire process
We now focus our attention on one of the steps with the larger impact on both prediction performance and execution time: retrieval. Indeed, Fig. 4 shows how re-ranking with exact matching outperforms the approximate matching-base MinHash.
Figure 8. Relationship between containment and performance. (a) Average top-200 containment with different retrieval methods. (b) Regression plot relating the prediction performance with the Jaccard Containment; each dot represents an experimental run for each base table.
Fig. 8(a) shows the containment results for the tables obtained by the different retrieval methods. We report top-200 containment to highlight how, even when retrieving a quite large number of candidates, the average containment of MinHash is very low compared to the other methods: for larger datasets, MinHash returns thousands of candidates of which only a fraction is selected; the lack of an internal ranking means that high-containment candidates are likely to be missed (subsection 4.1). In fact, MinHash has a threshold that can be tweaked to reduce the number of candidates that are retrieved, however we have observed that recall drops sharply at high thresholds. Hybrid Minhash has limited success in mitigating the problem.  [^4] The performance of MinHash on Binary is an artefact: the method could only retrieve about 6 candidates on average compared to the 30 found by the other two methods, thus values are averaged across fewer candidates with higher containment. We were unable to obtain the containment results for Starmie on YADL 50k and Open Data US.
We now focus on the best performing configuration (Exact Matching, Stepwise Greedy Join, Catboost). Fig. 8(b) represents the relationship between containment and prediction performance. The plotted line highlights the clear upwards trend in the prediction performance that follows the increase in containment. In other words, higher prediction results occur more frequently when the containment is higher. When the join is exact (like in our scenario), having a high Jaccard containment leads to better results because left joins with high overlap augment a large fraction of the base table, whereas low overlap joins produce features that contain mostly empty values.
While Hybrid MinHash appears to be faster than Exact Matching in the pipeline, it is important to note that re-ranking candidates incurs a non-negligible additional cost (Fig. 4, Table 5).
5.3. Smart aggregation brings marginal benefits
While containment impacts heavily both prediction performance and execution time, aggregation has a similar impact on the execution time without the same degree of improvement for the prediction performance (Fig. 5c). Aggregating values always leads to a loss of information: in exchange for a larger cost, complex aggregation methods that preserve more information (DFS) or that replace values with better representatives of a group (mean) should lead to better prediction performance.
However, this is not what we observe: the improvement in prediction performance of DFS is small in comparison with its large computational cost. This problem is exacerbated by the fact that aggregation must be performed whenever a join is executed at any point in the pipeline, including joins executed during join selection. The result is a compounding slow-down of the entire pipeline.
The similar performance of “first” and “mean” can be explained by the execution of joins on key-like columns so that values in these columns tend to be mostly distinct, thus producing the same aggregated value for both “mean” and “first”. The slight difference in performance and runtime between the two methods may be attributed to the large fraction of missing values in some of the data lakes: the “mean” method is biased towards selecting the most frequent value, which may be a null. When that happens, more values than needed become nulls, reducing the amount of features and introducing a slight speedup in the overall training.
The poor performance of DFS is also explained by the fact that we are not fully exploiting its capabilities. We consider only join depth-1 chains: at each aggregation step, we join the base table with an additional table, rather than leveraging the recursive generation of features provided by DFS. As a result, DFS is not as effective at generating features as it would be with deeper join paths (Cvetkov-Iliev et al., 2023) and may benefit from integrating join path discovery systems (Deng et al., 2017a).
5.4. Tradeoff between analytic performance and compute cost
Figure 9. Breakdown of where time is spent for each selector. On the right we report the average run time of each selector.
Execution time
Fig. 9 reports a breakdown of where the time is spent by different join estimation methods with the corresponding average total time on the right. Prepare tracks the time spent building data structures and loading data, including loading candidate tables in memory and building the containment ranking. Train(model) and Predict(model) track the time spent inside the ML model for training and prediction, respectively. Finally, Train(join) and Predict(join) track the time spent executing a merge operation, combining join and aggregation. Results are aggregated over all experiments with “first” as aggregation. The time distribution follows our expectations: most of the time is spent fitting models, with time spent joining (and aggregating) in second place. Highest Containment Join spends a relatively long time in the “Prepare” step due to the need to re-rank candidates before joining: since the join and train steps involve only one table, they are faster in comparison. Stepwise Greedy Join has a similar re-ranking step, however training the models in each iteration dominates the other steps. Full Join involves merging all candidates at the same time, then training a single model on the result: this explains how the fraction of time spent joining is comparatively larger than in other methods. Top-$1$ Full Join involves merging a single table without preparation, which makes it very fast.
Table 5. Average time (in s) spent building the index and querying all the base tables.
Retrieval M. Time Build Time Query
Exact Matching 1982.1 0.48
MinHash 3557.8 141.9
H. MinHash 1715.3
Table 5 reports the average time across all data lakes and base tables to build data structures and query candidates. Hybrid MinHash and MinHash have the same time as they share the data structure.
Overall, preparing a single column for querying with Exact Matching requires about 300 seconds on average (when parallelizing over 32 threads), and this cost increases linearly with the number of columns to query, the size of the data lake and the size of the base table. Conversely, building the MinHash index takes much longer than a single column, but it is a one-time operation to be executed on the lake and scales with the size of the data lake itself. Hybrid MinHash reuses the basic MinHash index and has therefore the same creation time; query time is much longer, however, owing to the need to re-rank the query results.
Overall, the time spent performing retrieval depends mostly on the number of performed query operations, rather than on the indexing operation. In our experiments, the time to measure the exact containment on the YADL Base data lake overtook the time for building the MinHash index after about 5 queries. This is also a consequence of how MinHash is designed specifically to minimize query time (Zhu et al., 2023, 2016).
Figure 10. Evolution of the runtime as a function of the number of query columns for the different retrieval methods.
The values used in Fig. 10 are obtained assuming that query retrieval time increases linearly with the size of a table and that the cost of creating the MinHash index is fixed for the data lake at hand. It shows how the fixed cost involved in building the MinHash index pays off after as few as six queries on average thanks to the fast query time. Hybrid MinHash requires more queries to break even due to its slower query time, yet it remains faster than Exact Matching when the number of query columns is larger. While the assumptions may not hold in general, the plot gives a reasonable estimate of the break-even points.
Although dependent on implementation, another factor that should be considered is the size on disk of the indices: the MinHash index occupies a much larger space on disk compared to the data required to hold the Exact matching ranking. Starmie needs to store model checkpoints, which occupy roughly 500MB regardless of the size of the base data lake (Table 6).
Concluding, the choice of the retrieval method falls on the user. Exhaustive computation of the containment is a net gain in performance at the expense of an execution time that increases quickly as the number of columns to query increases. This may not be a problem if the user is aware of which columns should be queried; if, instead, the user is trying to conduct an exhaustive search over all columns, a method such as MinHash should be favored. These observations are consistent with (Zhu et al., 2019). Finally, in scenarios where the query table and the data lake do not change, query results can computed offline and reused; in these scenarios, the additional cost of Exact Matching would be less problematic. In situations where the data lake tends to evolve over time, methods that support updating the index such as (Fernandez et al., 2019) or (Fernandez et al., 2018) should be considered.
Table 6. Recorded peak RAM usage for building and querying the different retrieval methods. All sizes are reported in MB. Runs that failed are marked as “NA”.
Case
Data Lake
Version
Exact
Matching
Minhash
Minhash
Hybrid
Starmie
Disk Binary 0.0 1.4 1.4 501.1
Open Data US 12.7 434 434 NA
YADL 10k 21.3 515 515 522.3
YADL 50k 28.2 2480 2480 NA
YADL base 21.3 462 462 522.4
RAM Build Binary 1196 226 226 19424
Open Data US 5466 3288 3286 NA
YADL 10k 688 3312 3290 87716
YADL 50k 1259 13697 13697 NA
YADL base 1061 2810 2810 38683
RAM Query Binary 255 216 216 4567
Open Data US 384 7789 10472 NA
YADL 10k 388 9138 12679 134869
YADL 50k 781 37332 37327 NA
YADL base 412 6784 8541 145437
5.5. Take-away messages
We summarize the main take-away messages from the experiments.
  1. (1)
    Retrieving the correct join candidates is the most important task of the pipeline; good retrieval yields higher quality candidates, which improves performance in later steps (Figures 4, 8a, 8b).
  2. (2)
    Join Selection cannot compensate for poor retrieval performance (Figure 5a). Joining all candidates from the retrieval step is the best trade-off between resource utilization and downstream performance.
  3. (3)
    Complex aggregation methods are much slower than simpler ones and do not result in commensurate gains in prediction performance (Figure 5b). Furthermore, complex methods scale worse with multi-table selectors.
  4. (4)
    Regardless of the ML model, execution time is dominated by the training when the aggregation is simple, while the inverse is true with complex aggregation functions (Figure 9).
  5. (5)
    Hybrid retrieval methods mitigate some of the drawbacks of stochastic methods (Figure 8) and scale well with many query columns (Figure 10).
  6. (6)
    Joining candidates brings benefits over working with the base table (Fig. 7). The characteristics of a data lake (Table 3, Figure 2) affect many facets of the results (Figures 7, 8).
6. Conclusion
In this work, we build a data lake based on real data contained in a KB to use as test bed for evaluating methods to augment user-provided tables. We implement an easily-extendable augmentation pipeline to test the different steps of the augmentation procedure and alternative algorithms for each of their tasks. Our results uncover a number of observations that we believe could help with directing research on this subject.
Future work. There are various interesting possible extensions.
More Lakes
Experiments show that YADL is representative of real data lakes. However, it mostly features categorical data. This characteristic partially explains the poor performance of DFS in the aggregation task. Future work could include deploying more YADL variants, possibly with domain-specific data lakes (Shuja et al., 2021).
More Tables
The tables we studied are small enough for Full Join and Stepwise Greedy Join to be viable. Larger tables would stress the pipeline and probe methods scalability even more. While we report more base tables than previous work, yet a larger pool would be valuable. An orthogonal problem is evaluating ethical issues in the choice of base tables and lake tables (Nargesian et al., 2022).
More Methods
We do not try to be exhaustive over the possible methods as we believe the implemented ones are representatives of the recent literature for this problem. However, there are more design choices that have an impact on the performance. We report here a few examples of methods that could be deployed and tested in our pipeline. (1) The selectors in our pipeline join a full table at each iteration. This adds multiple features without executing the same join multiple times, but it has the risk of adding noise by introducing irrelevant columns. A possible solution is executing joins that add one column at a time (Galhotra et al., 2023). Alternatively, methods such as (Esmailoghli et al., 2021) could be used to select retrieval candidates with high correlation. (2) It has been shown that data lake table profiling can bring in benefit in selecting joins and augmenting data (Galhotra et al., 2023; Flores et al., 2021). Profiling helps reducing the search space and directing retrieval methods towards meaningful candidates. (3) We report on the scenario where it is possible to perform joins with equality. However, this is not possible in general because of typos, different formats, and different granularity. Similarity joins would help with this problem (Jiang et al., 2014; Dong et al., 2023; Deng et al., 2017b; Mundra et al., 2023) and add another dimension to the analysis.
Chain of Joins
To limit the search space, we keep our chain of joins for augmentation limited to one. However, some of the methods have considered chains of join for augmentation (Galhotra et al., 2023). Enabling join chains would also make clear the benefit of recursive methods such as DFS (Cvetkov-Iliev et al., 2023).
Augmentation Beyond Joins
Another popular augmentation technique is “table unionability”: tables are considered unionable if they have the same structure, i.e., they have the same number of columns with compatible data types  (Bogatu et al., 2020; Khatiwada et al., 2023a). Also, table discovery over data lakes includes other approaches that are not driven by the base table and joins (Fan et al., 2023a), for example, are driven by user keywords (Halevy et al., 2016) or by the embedding of the table (Deng et al., 2022; Hulsebos et al., 2023b).
Final words. Our study can be very useful to the practitioner as it reveals a bag of simple tricks to automate supervised learning on a data lake with fairly good statistical performance and yet good scaling with the size of the data lake.
We acknowledge the invaluable insight and expertise provided by Léo Dreyfus-Schmidt and Du Phan (formerly from Dataiku), without whom the foundations of this work would not have existed.
As YADL is based on YAGO 3, we acknowledge the work made by its authors to prepare the original knowledge base, as well as their efforts into manually evaluating it.